OpenStudy (anonymous):

what is the slope and y intercept of 1/3y-1/2x>2

OpenStudy (anonymous):

\[1/3y-1/2x>2\] This is an inequality rather than just a line, but you can treat it as a regular linear equation of form y=mx+b to find the slope and y-intercept. First add 1/2x to each side of the inequality to isolate the y variable, obtaining \[1/3y>1/2x+2\] Then, multiply both sides by 3 to obtain \[3*1/3y>3*(1/2x+2)\] or \[y>3/2x+6\] This is form y>mx+b, where slope is m and y-intercept is b. So the slope is 3/2 and the y-intercept is 6.

OpenStudy (anonymous):

on a graph where would the shaded region be? and would it be a solid or dotted line?

OpenStudy (anonymous):

Since it is greater than, it would be above the line (think of it as all the lines above the line y=3/2x+6, but not including y=3/2x+6). And the line would be dotted because it is not greater than or equal to, just greater than.

OpenStudy (anonymous):

ok thank you